OpenStudy (shadowcain):

Mrs. Moore's class is conducting an experiment to find the probability of pulling certain colors from a bag of 25 marbles. If 6 are purple, 5 are yellow, 3 are green, and the rest are black, what is the probability of drawing 2 black if the marbles are not replaced after they are picked? Show answer as a percent rounded to nearest hundredth, if necessary. A. 17.6% B. 18.33% C. 19.36% D. 20.17%

OpenStudy (shadowcain):

Possible outcomes 1. 2 purple 2. 2 yellow 3. 2 green 4. 2 black 5. 1 p, 1 y 6. 1 p, 1 g 7. 1 p, 1 b 8. 1 y, 1 g 9. 1 y, 1 b 10. 1 g, 1 b

OpenStudy (shadowcain):

@Michele_Laino

OpenStudy (michele_laino):

we have: 25-(6+5+3)=11 black marbles so at first draw, the probability to get a black marble is: favorable outcomes/ possible outcomes= 11/25 such probability at the second draw, is: favorable outcomes/ possible outcomes= 10/24 since after the first draw, we have remained with 10 black marbles and 24 marbles in total Now the requested probability is given by the product of such probabilities

OpenStudy (shadowcain):

But how do I get the %